### v1

@@ 1 Introduction
In our prior work, we introduced the Hydrodynamic Interaction Graph Neural Network (HIGNN) [30], a new framework that harnesses the power of GNNs to enable fast simulation of particulate suspensions. This framework has subsequently been utilized to examine a variety of suspension drop behaviors, including shape deformation, disintegration, and coalescence [31]. The basic idea of HIGNN is to represent a suspension of particles as a graph, where the particles and two-body interactions are represented as vertices and edges, respectively. Besides, to incorporate HIs beyond two-body, i.e., $m$-body interactions with $m>2$, the graph is generalized to include higher-order connectivity. For example, a face connectivity between three vertices (particles) is introduced into the graph and used to describe three-body connections and thereby to capture the three-body HI effects. By introducing higher-order connectivity in the graph and utilizing corresponding convolutional operations, HIGNN enables accurately modeling many-body HIs between particles. As discussed, each connectivity within the graph, such as edges or faces, accounts for a corresponding type of $m$-body interactions (two-body or three-body, respectively). The GNN learns the mapping from these connectivities to their associated interactions. This direct correspondence between graph connectivity and physical interactions lends explainability to the HIGNN model, distinguishing it from “black-box” neural network models. Another notable advantage of HIGNN is its transferability across different systems. Similar to SD, it directly models the mobility tensor, and hence can be applied to simulating suspensions of particles subject to any type and magnitudes of external forces. The dependence of each $m$-body (e.g., two-body or three-body) interactions on the relative positions of associated particles remains invariant, irrespective of the total number or concentration of particles in the suspension. Therefore, once the HIGNN is trained for a domain, it can be applied to simulating suspensions of varying numbers of particles in the same domain. Furthermore, training HIGNN requires minimal computational resources: if achieving the desired accuracy necessitates preserving $m$-body HIs, the training process only requires data for up to $m$ particles, sampled across a range of spatial configurations within the domain of interest. HIGNN offers the potential for significant computational speedup by requiring only a forward pass through its neural networks at each time step, unlike traditional methodologies discussed above [32, 15, 16, 18, 11, 24] that typically require solving large-scale linear systems. However, HIGNN’s overall prediction cost scales quadratically, $O(N^{2})$, exceeding the $O(N\log N)$ scaling of ASD [24]. As a result, while HIGNN outperforms ASD for small- to moderate-scale problems ($N\leq 10^{4}$) [30, 31], its cost approaches or surpasses ASD for very large-scale simulations. This $O(N^{2})$ complexity stems from the slow decay of two-body HIs, necessitating edges between all particle pairs and a fully-connected graph representation of the particle system. This results in an $O(N^{2})$ time complexity for the matrix-vector multiplication during inference.
@@
To ensure consistent efficiency superiority across all scales, in the present work, we aim to reduce the prediction cost scaling of HIGNN to quasi-linear, i.e., $O(N\log N)$. To achieve this, we propose integrating the hierarchical matrix ($\mathcal{H}$-matrix) techniques [33, 34] into the HIGNN framework. The core principle of $\mathcal{H}$-matrix involves partitioning a matrix into blocks. The partitioned blocks meeting the admissibility condition [34] can then be approximated by low-rank matrices [35, 36]. This low-rank approximation can significantly reduce the cost of matrix-vector multiplication. Constructing an $\mathcal{H}$-matrix requires $O(N\log N)$ operations, while matrix-vector multiplication with an $\mathcal{H}$-matrix typically exhibits $O(N)$ complexity. Therefore, integrating $\mathcal{H}$-matrix addresses our need to reduce HIGNN’s computational complexity to quasi-linear. Moreover, as shown later in our numerical tests, the actual computing time associated with $\mathcal{H}$-matrix remains low. Consequently, even for simulations involving very large-scale particulate suspensions (e.g., containing more than millions of particles), computational costs related to hardware resources and wall time can be maintained at a low level. It is worth noting that our approach differs from previous work incorporating hierarchical structures into GNNs, such as Multipole Graph Neural Operator [37] and Hierarchical Message-Passing Graph Neural Networks [38]. These prior methods formulate the entire model structure as a $\mathcal{H}$-matrix and learn the message passing between different graph levels (i.e. low-rank approximation in hierarchical matrices) from data. Thus, these approaches require training datasets with large numbers of particles, which is not feasible in our context due to the huge cost of data generation. Instead, we maintain HIGNN’s advantage of training on data with small numbers of particles, while only applying $\mathcal{H}$-matrix techniques during the inference stage to build our $\mathcal{H}$-HIGNN framework.
@@
Integrating $\mathcal{H}$-matrices into the HIGNN framework requires a concerted technical effort, involving several key components. First, recognizing the inherent geometric structure of the particle system, we performed cluster-tree block partitioning [39] by using the geometric information embedded in the graph. Second, the admissibility is determined based on the relative distance between clusters and their respective sizes. Third, by leveraging this admissibility condition and the asymptotic smoothness of two-body HIs, we implemented an adaptive cross approximation (ACA) algorithm [36] to achieve effective low-rank approximations for a significant portion of partitioned blocks in the mobility matrix. Fourth, unlike conventional $\mathcal{H}$-matrices, where entries are typically scalars, each element in our mobility matrix is a $3\times 3$ submatrix. This necessitated the development of specialized matrix operations and data structures, especially in the code implementation, to efficiently handle these submatrices within the $\mathcal{H}$-matrix framework, ensuring compatibility and performance. Finally, through careful optimization of GPU latency, workload distribution, and inter-GPU communication, coupled with the Kokkos programming model [40], our implementation achieves a computer wall-time that closely approximates the theoretical $O(N\log N)$ scaling and yields a near-ideal strong scalability for parallel efficiency. Such $\mathcal{H}$-matrix accelerated HIGNN is designated as $\mathcal{H}$-HIGNN. Through a comprehensive evaluation, we assess the accuracy of $\mathcal{H}$-HIGNN against the original HIGNN, and validate its computational complexity and scalability. Furthermore, we demonstrate its efficiency in simulating practically large-scale particle suspensions. By incorporating elastic bonding and bending forces between particles, $\mathcal{H}$-HIGNN can also be applied to simulating large-scale suspensions of flexible filaments.
@@ 2.1 Problem setup
[…]
@@ 2.2 Theoretical foundation
[…]
@@ 2.3 Framework structure
In addition, HIGNN constructs two Multilayer Perceptron (MLP) neural networks $\mathbf{h}_{\boldsymbol{\Theta}_{2}}$ and $\mathbf{g}_{\boldsymbol{\Theta}_{3}}$ on edge and face connections as the surrogates of $\boldsymbol{\alpha}_{2}$ and $\boldsymbol{\alpha}_{3}$, respectively, with $\boldsymbol{\Theta}_{2}$ and $\boldsymbol{\Theta}_{3}$ the network parameters. To this end, the information on all edges and faces associated with the target vertex are aggregated by edge convolutional (EdgeConv) [43] and face convolutional (FaceConv) [30] operations, respectively. The main components of the HIGNN framework are summarized in FIG. 1.
@@
(b) The surrogate model of $\boldsymbol{\alpha}_{3}(\mathbf{X}_{j},\mathbf{X}_{k},\mathbf{X}_{i})$ is the MLP neural network $\mathbf{g}_{\boldsymbol{\Theta}_{3}}$ obtained through FaceConv.
@@ 2.4 Training
[…]
@@
In the generation of data involving two particles, the maximum distance considered between the two particles was set sufficiently large (e.g., 1000) so as to encompass the range of distances likely to be encountered in subsequent tests and simulations. All numerical values are non-dimensional, with the radius of particles normalized to 1. For the data involving three particles, used exclusively to train the three-body HIs, the maximum interparticle distance was set to the cutoff distance, $R_{\text{cut}}$. Our prior studies [30] have determined $R_{\text{cut}}=5.0$, and hence this cutoff distance is used to define face connectivity in the graph. For each dataset, particle positions were randomly generated within the established distance constraints, and each particle was assigned a unit force randomly sampled from the set: $\mathbf{F}=[1,0,0]^{T}$, $[0,1,0]^{T}$, or $[0,0,1]^{T}$. For the purposes of the present study, 100,000 distinct configurations of two particles and 50,000 distinct configurations of three particles were generated in an unbounded domain. For each configuration, a numerical solver can be invoked to calculate particle velocities for a single time step, given the particle positions and assigned forces. Due to the code accessibility, we used SD [44]. Each set of particle positions, applied forces, and resulting velocities constitutes a single data point to be utilized in the subsequent training and validation processes. The generation of each data point only requires a wall time of 0.014 seconds on a workstation equipped with an Intel(R) Xeon(R) E5-2698 v4 CPU @ 2.20GHz.
@@ 2.5 Prediction: transferability and cost
To predict the time evolution of particles, each particle’s position is updated from its velocity by numerically integrating $\dot{\textbf{X}}=\mathbf{U}$ over time. At each time step, the trained HIGNN is called to predict all particles’ velocities, using the graph $\boldsymbol{\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{F})}$ , which fully characterizes the particle system, as input. This allows us to effectively simulate the dynamics of particles in suspension.
HIGNN can potentially achieve significant computational speedup by requiring only a forward pass through its MLP neural networks at each time step, in contrast to traditional simulation methodologies [32, 15, 16, 18, 11, 24], which usually require solving a large-scale linear system at each time step. The computational complexity of HIGNN is directly related to the number of edges and faces in the graph, as neural network inferences are performed on each edge and face. As discussed above, the two-body HI not only dominates the short-range lubrication effect but also decays very slowly ($\sim O(r^{-1})$) in long range. Thus, for each particle, we need to consider all other particles as its neighbors for the calculation of $\boldsymbol{\alpha}^{(t)}_{2}$. In HIGNN, this translates to the construction of edges between any two particles, leading to a fully-connected graph characterized by $N(N-1)$ edges. As for the faces, which characterize three-body connections, the faster decaying rate of three-body HIs allows for the implementation of a finite cutoff distance ($R_{\text{cut}}$), thereby limiting the number of faces to $O(N)$. Therefore, the overall computational complexity (prediction cost) of HIGNN scales as $O(N^{2})$ with $N$ the total number of particles, which is greater than the scaling $O(N\log N)$ of ASD [24]. This implies that while HIGNN’s computational cost is significantly lower than ASD for small- to moderate-scale problems (as shown in our previous work [30, 31]), for simulations of very large-scale particulate suspensions, HIGNN’s cost can approach or surpass that of ASD. Therefore, to achieve a consistent superiority on computational efficiency, in the present work we aim to reduce the scaling of the prediction cost of HIGNN down to quasi-linear, i.e., ${O}(N\log N)$. By such, even for simulating very large-scale particulate suspensions (e.g., containing more than millions of particles), the computational cost can still be maintained low. For achieving that, we propose to integrate the $\mathcal{H}$-matrix techniques [33, 34] into the HIGNN framework. The detailed methodology is introduced in the next section.
3 Hierarchical Matrix ($\mathcal{H}$-Matrix)--Accelerated HIGNN
@@
where $\mathbf{M}^{(1)}$, $\mathbf{M}^{(s,2)}$ and $\mathbf{M}^{(s,3)}$ are all diagonal matrices with each element (a $3\times 3$ submatrix) calculated from $\boldsymbol{\alpha}_{1}$, $\boldsymbol{\alpha}^{(s)}_{2}$ and $\boldsymbol{\alpha}^{(s)}_{3}$, respectively; the matrix $\mathbf{M}^{(t,2)}$ is constructed from the two-body interactions $\boldsymbol{\alpha}_{2}^{(t)}$, and $\mathbf{M}^{(t,3)}$ from the three-body interactions $\boldsymbol{\alpha}_{3}^{(t)}$. Following the preceding analysis, the matrix-vector multiplications appearing in Eq. (6) typically exhibit $O(N)$ complexity due to the sparsity of the associated matrices, but with the exception of the multiplication $\mathbf{M}^{(t,2)}\cdot\mathbf{F}$, which has $O(N^{2})$ complexity because $\mathbf{M}^{(t,2)}$ is a dense matrix. Thus, we integrate the $\mathcal{H}$-matrix techniques [33, 34] to reduce the computational complexity for the multiplication $\mathbf{M}^{(t,2)}\cdot\mathbf{F}$ from $O(N^{2})$ to $O(N\log N)$. Note that $\mathbf{M}^{(t,2)}$ is in fact assembled by $N\times N$ submatrices, where each submatrix $\mathbf{M}^{(t,2)}_{ij}\in\mathbb{R}^{3\times 3}$ corresponds to $\boldsymbol{\alpha}_{2}^{(t)}(\mathbf{X}_{i},\mathbf{X}_{j})$. Following the convention of $\mathcal{H}$-matrix, it is more convenient to let the matrix’s dimension equal to the number of particles. Hence, we denote each submatrix $\mathbf{M}^{(t,2)}_{ij}$ as an element or entry of the matrix, and thereby $\mathbf{M}^{(t,2)}$ is referred to as an $N\times N$ matrix. [^1] For the subsequent discussion, it should be noted that when performing any operation on elements, each element corresponds to a $3\times 3$ submatrix, rather than a scalar value.
@@
The basic idea of $\mathcal{H}$-matrix is to partition $\mathbf{M}^{(t,2)}$ into blocks based on geometric information. The blocks satisfying the admissibility condition [34], referred to as the admissible blocks, can then be approximated by low-rank matrices [35, 36]. Intuitively, an admissible block can be regarded as a block where the associated vertices exhibit relatively longer edge connections. A more rigorous mathematical definition is given in §3.1. After constructing the low-rank approximation, the complexity of the matrix-vector multiplication would be effectively reduced. For example, considering two subsets of vertices (or particles), one with $c$ particles, and the other with $q$ particles, the block representing the two-body interactions between these two sets of particles is denoted as $\mathbf{G}\in\mathbb{R}^{c\times q}$. Its $l$-rank approximation can be expressed as: $\mathbf{G}\approx\mathbf{C}\mathbf{Q}$, where $\mathbf{C}\in\mathbb{R}^{c\times l}$ and $\mathbf{Q}\in\mathbb{R}^{l\times q}$. When multiplying with a vector $\mathbf{b}\in\mathbb{R}^{q}$, we can first calculate the product $\mathbf{Q}\mathbf{b}$ and then multiply the resultant vector with $\mathbf{C}$. By such, the number of operations required can be reduced from ${O}(cq)$ to ${O}((c+q)l)$. Since the function $\boldsymbol{\alpha}_{2}^{(t)}$ is asymptotically smooth, $\mathbf{G}$ can be well approximated by the $l$-rank approximation with $l\ll c$ and $l\ll q$, as long as $\mathbf{G}$ satisfies the admissibility condition [34, 35, 36]. Thus, the cost of the matrix-vector multiplication can be significantly reduced by the low-rank approximation. Given a configuration $\mathbf{X}$ of $N$ particles, we can employ the cluster-tree partitioning algorithm [39] to partition the associated matrix $\mathbf{M}^{(t,2)}$ into blocks such that each block satisfying the admissibility condition can be well approximated by a low-rank approximation. The resulting matrix, composed largely of low-rank blocks, is the so-called $\mathcal{H}$-matrix.
@@
In this work, we employ the adaptive cross approximation (ACA) [36] to find a low-rank approximation, thereby avoiding the computation of all entries of the original matrix. As a result, constructing $\mathcal{H}$-matrix requires ${O}(N\log N)$ operations, whereas the matrix-vector multiplication with an $\mathcal{H}$-matrix exhibits ${O}(N)$ complexity. More details about the cluster-tree block partitioning and ACA are provided as follows.

### v2

@@ 1 Introduction
In our prior work, we introduced the Hydrodynamic Interaction Graph Neural Network (HIGNN) [30], a new framework that harnesses the power of GNNs to enable fast simulation of particulate suspensions. This framework has subsequently been utilized to examine a variety of suspension drop behaviors, including shape deformation, disintegration, and coalescence [31]. The basic idea of HIGNN is to represent a suspension of particles as a graph, where the particles and two-body interactions are represented as vertices and edges, respectively. Besides, to incorporate HIs beyond two-body, i.e., $m$-body interactions with $m>2$, the graph is generalized to include higher-order connectivity. For example, a 3-clique connectivity between three vertices (particles) is introduced into the graph and used to describe three-body connections and thereby to capture the three-body HI effects. By introducing higher-order connectivity in the graph and utilizing corresponding convolutional operations, HIGNN enables accurately modeling many-body HIs between particles. As discussed, each connectivity within the graph, such as edge or 3-clique, accounts for a corresponding type of $m$-body interactions (two-body or three-body, respectively). The GNN learns the mapping from these connectivities to their associated interactions. This direct correspondence between graph connectivity and physical interactions lends explainability to the HIGNN model, distinguishing it from “black-box” neural network models. Another notable advantage of HIGNN is its transferability across different systems. Similar to SD, it directly models the mobility tensor, and hence can be applied to simulating suspensions of particles subject to any type and magnitudes of external forces. The dependence of each $m$-body (e.g., two-body or three-body) interactions on the relative positions of associated particles remains invariant, irrespective of the total number or concentration of particles in the suspension. Therefore, once the HIGNN is trained for a domain, it can be applied to simulating suspensions of varying numbers of particles in the same domain. Furthermore, training HIGNN requires minimal computational resources: if achieving the desired accuracy necessitates preserving $m$-body HIs, the training process only requires data for up to $m$ particles, sampled across a range of spatial configurations within the domain of interest. HIGNN offers the potential for significant computational speedup by requiring only a forward pass through its neural networks at each time step, unlike traditional methodologies discussed above [32, 15, 16, 18, 11, 24] that typically require solving large-scale linear systems. However, HIGNN’s overall prediction cost scales quadratically, $O(N^{2})$, exceeding the $O(N\log N)$ scaling of ASD [24]. As a result, while HIGNN outperforms ASD for small- to moderate-scale problems ($N\leq 10^{4}$) [30, 31], its cost approaches or surpasses ASD for very large-scale simulations. This $O(N^{2})$ complexity stems from the slow decay of two-body HIs, necessitating edges between all particle pairs and a fully-connected graph representation of the particle system. This results in an $O(N^{2})$ time complexity for the matrix-vector multiplication during inference.
@@
To ensure consistent efficiency superiority across all scales, in the present work, we aim to reduce the prediction cost scaling of HIGNN to quasi-linear, i.e., $O(N\log N)$. To achieve this, we propose integrating the hierarchical matrix ($\mathcal{H}$-matrix) techniques [33, 34] into the HIGNN framework. The core principle of $\mathcal{H}$-matrix involves partitioning a matrix into blocks. The partitioned blocks meeting the admissibility condition [34] can then be approximated by low-rank matrices [35, 36]. This low-rank approximation can significantly reduce the cost of matrix-vector multiplication. Constructing an $\mathcal{H}$-matrix requires $O(N\log N)$ operations, while matrix-vector multiplication with an $\mathcal{H}$-matrix typically exhibits $O(N)$ complexity. Therefore, integrating $\mathcal{H}$-matrix addresses our need to reduce HIGNN’s computational complexity to quasi-linear. Moreover, as shown later in our numerical tests, the actual computing time associated with $\mathcal{H}$-matrix remains low. Consequently, even for simulations involving very large-scale particulate suspensions (e.g., containing more than millions of particles), computational costs related to hardware resources and wall time can be maintained at a low level. It is worth noting that our approach differs from previous work incorporating hierarchical structures into GNNs, such as Multipole Graph Neural Operator [37] and Hierarchical Message-Passing Graph Neural Networks [38]. These prior methods formulate the entire model structure as a $\mathcal{H}$-matrix and learn the message passing between different graph levels (i.e. low-rank approximation in $\mathcal{H}$-matrices) from data. Thus, these approaches require training datasets containing large numbers of particles, which is not feasible in our context due to the high cost of data generation. Instead, we preserve HIGNN’s advantage of training on data with small numbers of particles, while only applying $\mathcal{H}$-matrix techniques during the inference stage to build our $\mathcal{H}$-HIGNN framework.
@@
Integrating $\mathcal{H}$-matrices into the HIGNN framework requires a concerted technical effort, involving several key components. First, recognizing the inherent geometric structure of the particle system, we performed cluster-tree block partitioning [39] by using the geometric information embedded in the graph. Second, the admissibility is determined based on the relative distance between clusters and their respective sizes. Third, by leveraging this admissibility condition and the asymptotic smoothness of two-body HIs, we implemented an adaptive cross approximation (ACA) algorithm [36] to achieve effective low-rank approximations for a significant portion of partitioned blocks in the mobility matrix. Fourth, unlike conventional $\mathcal{H}$-matrices, where entries are typically scalars, each element in our mobility matrix is a $3\times 3$ submatrix. This necessitated the development of specialized matrix operations and data structures in both the ACA algorithm and code implementation to efficiently handle these submatrices within the $\mathcal{H}$-matrix framework, ensuring compatibility and performance. Finally, through careful optimization of GPU latency, workload distribution, and inter-GPU communication, coupled with the Kokkos programming model [40], our implementation achieves a computer wall-time that closely approximates the theoretical $O(N\log N)$ scaling and yields a near-ideal strong scalability for parallel efficiency. Such $\mathcal{H}$-matrix accelerated HIGNN is designated as $\mathcal{H}$-HIGNN. Through a comprehensive evaluation, we assess the accuracy of $\mathcal{H}$-HIGNN against the original HIGNN, and validate its computational complexity and scalability. Furthermore, we demonstrate its efficiency in simulating practically large-scale particle suspensions. By incorporating elastic bonding and bending forces between particles, $\mathcal{H}$-HIGNN can also be applied to simulating large-scale suspensions of flexible filaments.
@@ 2.1 Problem setup
Two types of traditional approaches exist for obtaining the mobility matrix $\textbf{M}(\textbf{X})$: directly solving the Stokes equations with prescribed boundary conditions on the particles’ surfaces [12, 16, 17, 18]; and approximating $\textbf{M}(\textbf{X})$ through multipole expansions truncated at the stresslet level, as in SD and its variants [11, 22, 24, 41]. However, both methods are computationally demanding, and the computed $\textbf{M}(\textbf{X})$ lacks transferability to other suspensions even with the same type of particles but different particle numbers or concentrations. To mitigate these limitations, we have proposed the HIGNN framework [30] that enables fast simulations and transferability across systems with different numbers of particles in the same domain.
@@ 2.2 Theoretical foundation
Based on the many-body nature of HI [42] and our analysis [30], the mobility tensor $\textbf{M}(\textbf{X})$ can be expressed as an expansion of additive terms arising from single-body and different $m$-body contributions, e.g., two-body, three-body, etc. From theoretical analysis on the decaying orders of different $m$-body contributions with respect to the characteristic particles’ separation distance (denoted as $r$) [42, 30], we know that for greater $m$, the corresponding $m$-body contribution decays at a faster rate in terms of $r$. In the original HIGNN framework [30], the $m$-body contributions with $m\geq 4$ are truncated, retaining only two- and three-body interactions in HI. While used for proof of concept, this approximation is not inherent to the framework and has been validated for suspensions that are not too concentrated, achieving the desired level of accuracy [30, 31]. Henceforth, the velocity of each particle can be approximated as:
@@ 2.3 Framework structure
In addition, HIGNN constructs two Multilayer Perceptron (MLP) neural networks $\mathbf{h}_{\boldsymbol{\Theta}_{2}}$ and $\mathbf{g}_{\boldsymbol{\Theta}_{3}}$ on edge and 3-clique connections as the surrogates of $\boldsymbol{\alpha}_{2}$ and $\boldsymbol{\alpha}_{3}$, respectively, with $\boldsymbol{\Theta}_{2}$ and $\boldsymbol{\Theta}_{3}$ the network parameters. To this end, the information on all edges and 3-cliques associated with the target vertex are aggregated by edge convolutional (EdgeConv) [43] and 3-clique convolutional (3-CliqueConv, referred to as FaceConv in [30, 31]) operations, respectively. The main components of the HIGNN framework are summarized in FIG. 1.
@@
(b) The surrogate model of $\boldsymbol{\alpha}_{3}(\mathbf{X}_{j},\mathbf{X}_{k},\mathbf{X}_{i})$ is the MLP neural network $\mathbf{g}_{\boldsymbol{\Theta}_{3}}$ obtained through 3-CliqueConv (referred to as FaceConv in the schematic following [30]).
@@ 2.4 Training
The training of HIGNN requires minimal computational resources, as discussed below. Training HIGNN to determine $\mathbf{h}_{\boldsymbol{\Theta}_{2}}$ and $\mathbf{g}_{\boldsymbol{\Theta}_{3}}$, for use in Eq. (3), requires only data of two or three particles sampled with different spatial configurations in the domain of interest. In general, if the attainment of desired accuracy necessitates the retention of $m$-body contributions in the mobility matrix $\textbf{M}(\textbf{X})$, the training process requires data comprising up to $m$ particles, sampled across a range of spatial configurations within the domain. For bounded domains, we need to train another neural network as the surrogate for $\boldsymbol{\alpha}_{1}(\mathbf{X}_{i})$, using the data of a single particle at different spatial positions in the domain. However, for unbounded and periodic domains, $\boldsymbol{\alpha}_{1}$ admits an analytical solution. Also, for particles in unbounded or periodic domains, HIs only depend on their relative positions or interparticle distances. Thus, normalizing the neural network outputs by the corresponding interparticle distances can improve the training performance. Eq. (3) can then be rewritten as:
@@
In the generation of data involving two particles, the maximum distance considered between the two particles was set sufficiently large (e.g., 1000) so as to encompass the range of distances likely to be encountered in subsequent tests and simulations. All numerical values are non-dimensional, with the radius of particles normalized to 1. For the data involving three particles, used exclusively to train the three-body HIs, the maximum interparticle distance was set to the cutoff distance, $R_{\text{cut}}$. Our prior studies [30] have determined $R_{\text{cut}}=5.0$, and hence this cutoff distance is used to define 3-clique connectivity in the graph. For each dataset, particle positions were randomly generated within the established distance constraints, and each particle was assigned a unit force randomly sampled from the set: $\mathbf{F}=[1,0,0]^{T}$, $0,1,0]^{T}$, or $0,0,1]^{T}$. For the purposes of the present study, 100,000 distinct configurations of two particles and 50,000 distinct configurations of three particles were generated in an unbounded domain. For each configuration, a numerical solver can be invoked to calculate particle velocities for a single time step, given the particle positions and assigned forces. Due to the code accessibility, we used SD [44]. Each set of particle positions, applied forces, and resulting velocities constitutes a single data point to be utilized in the subsequent training and validation processes. The generation of each data point only requires a wall time of 0.014 seconds on a workstation equipped with an Intel(R) Xeon(R) E5-2698 v4 CPU @ 2.20GHz.
@@ 2.5 Prediction: transferability and cost
As HIGNN essentially models the mobility tensor, the trained HIGNN is transferable to predicting the velocities of particles (suspended in fluid) subject to any types and magnitudes of external forces. Moreover, all $\boldsymbol{\alpha}_{i}$ terms in $\textbf{M}(\textbf{X})$ are invariant regardless of the number of particles, and hence the trained surrogate models for $\boldsymbol{\alpha}_{i}$ in HIGNN are transferable across suspensions of different concentrations or numbers of particles in the same domain. Because of its inherent transferability, the trained HIGNN eliminates the need for retraining when simulating particulate suspensions of varying concentrations and particle numbers under arbitrary external forces.
@@
HIGNN can potentially achieve significant computational speedup by requiring only a forward pass through its MLP neural networks at each time step, in contrast to traditional simulation methodologies [32, 15, 16, 18, 11, 24], which usually require solving a large-scale linear system at each time step. The computational complexity of HIGNN is directly related to the number of edges and 3-cliques in the graph, as neural network inferences are performed on each edge and 3-clique. As discussed above, the two-body HI not only dominates the short-range lubrication effect but also decays very slowly ($\sim O(r^{-1})$) in long range. Thus, for each particle, we need to consider all other particles as its neighbors for the calculation of $\boldsymbol{\alpha}^{(t)}_{2}$. In HIGNN, this translates to the construction of edges between any two particles, leading to a fully-connected graph characterized by N(N1)N(N-1) edges. As for the 3-cliques, which characterize three-body connections, the faster decaying rate of three-body HIs allows for the implementation of a finite cutoff distance ($R_{\text{cut}}$), thereby limiting the number of 3-cliques to $O(N)$. Therefore, the overall computational complexity (prediction cost) of HIGNN scales as $O(N^{2})$ with $N$ the total number of particles, which is greater than the scaling $O(N\log N)$ of ASD [24]. This implies that while HIGNN’s computational cost is significantly lower than ASD for small- to moderate-scale problems (as shown in our previous work [30, 31]), for simulations of very large-scale particulate suspensions, HIGNN’s cost can approach or surpass that of ASD. Therefore, to achieve a consistent superiority on computational efficiency, in the present work we aim to reduce the scaling of the prediction cost of HIGNN down to quasi-linear, i.e., ${O}(N\log N)$. By such, even for simulating very large-scale particulate suspensions (e.g., containing more than millions of particles), the computational cost can still be maintained low. For achieving that, we propose to integrate the $\mathcal{H}$-matrix techniques [33, 34] into the HIGNN framework. The detailed methodology is introduced in the next section.
3 Hierarchical Matrix ($\mathcal{H}$-Matrix)-Accelerated HIGNN
@@
where $\mathbf{M}^{(1)}$, $\mathbf{M}^{(s,2)}$ and $\mathbf{M}^{(s,3)}$ are all diagonal matrices with each element (a $3\times 3$ submatrix) calculated from $\boldsymbol{\alpha}_{1}$, $\boldsymbol{\alpha}^{(s)}_{2}$ and $\boldsymbol{\alpha}^{(s)}_{3}$, respectively; the matrix $\mathbf{M}^{(t,2)}$ is constructed from the two-body interactions $\boldsymbol{\alpha}_{2}^{(t)}$, and $\mathbf{M}^{(t,3)}$ from the three-body interactions $\boldsymbol{\alpha}_{3}^{(t)}$. Following the preceding analysis, the matrix-vector multiplications appearing in Eq. (6) typically exhibit $O(N)$ complexity due to the sparsity of the associated matrices, but with the exception of the multiplication $\mathbf{M}^{(t,2)}\cdot\mathbf{F}$, which has $O(N^{2})$ complexity because $\mathbf{M}^{(t,2)}$ is a dense matrix.
We determined that neither the Particle-Mesh Ewald (PME) method nor the Fast Multipole Method (FMM) were suitable for our problem. The PME method was originally designed for periodic boundary conditions [45] and achieves optimal performance in this context. While extensions were developed for non-periodic and semi-periodic systems [46, 47], these modifications could introduce additional computational overhead and errors. Thus, PME is not suitable for our problem scope, which aims to develop a method applicable to any type of boundary condition. Moreover, applying PME in our work would necessitate computing Fourier transforms of neural network-based interaction terms, introducing substantial additional computational expense. FMM typically requires analytical expansions, such as multipole, Legendre, or Taylor series [48, 49]. However, $\boldsymbol{\alpha}_{2}^{(t)}(\mathbf{X}_{i},\mathbf{X}_{j})$ is modeled by a neural network herein, and hence it is infeasible to represent $\boldsymbol{\alpha}_{2}^{(t)}$ using any analytical expansions. While it is possible to approximate $\boldsymbol{\alpha}_{2}^{(t)}$ with a sum of functions that have known multipole or other analytical expansions, this approach may compromise accuracy and incur additional computational cost due to the need to sample data from the neural network to construct the approximation. Variants like the kernel-independent FMM (KIFMM) [50] replace analytical expansions with interpolation (e.g., Chebyshev or polynomial), but its interpolation overhead can be also significant for neural network kernel evaluations. Finally, while FMM generally requires less setup time, $\mathcal{H}$-matrix tends to offer superior efficiency for matrix-vector multiplications and better parallel scalability [49, 51, 52], making it more suitable for the large-scale systems and numerous matrix-vector operations central to this work.
Thus, we integrate the $\mathcal{H}$-matrix techniques [33, 34] to reduce the computational complexity for the multiplication $\mathbf{M}^{(t,2)}\cdot\mathbf{F}$ from $O(N^{2})$ to $O(N\log N)$. Note that $\mathbf{M}^{(t,2)}$ is in fact assembled by $N\times N$ submatrices, where each submatrix $\mathbf{M}^{(t,2)}_{ij}\in\mathbb{R}^{3\times 3}$ corresponds to $\boldsymbol{\alpha}_{2}^{(t)}(\mathbf{X}_{i},\mathbf{X}_{j})$. Following the convention of $\mathcal{H}$-matrix, it is more convenient to let the matrix’s dimension equal to the number of particles. Hence, we denote each submatrix $\mathbf{M}^{(t,2)}_{ij}$ as an element or entry of the matrix, and thereby $\mathbf{M}^{(t,2)}$ is referred to as an $N\times N$ matrix. [^1] For the subsequent discussion, it should be noted that when performing any operation on elements, each element corresponds to a $3\times 3$ submatrix, rather than a scalar value.
@@
The basic idea of $\mathcal{H}$-matrix is to partition $\mathbf{M}^{(t,2)}$ into blocks based on geometric information. The blocks satisfying the admissibility condition [34], referred to as the admissible blocks, can then be approximated by low-rank matrices [35, 36]. Intuitively, an admissible block can be regarded as a block where the associated vertices exhibit relatively longer edge connections. A more rigorous mathematical definition is given in §3.1. After constructing the low-rank approximation, the complexity of the matrix-vector multiplication would be effectively reduced. For example, considering two subsets of vertices (or particles), one with $c$ particles, and the other with $q$ particles, the block representing the two-body interactions between these two sets of particles is denoted as $\mathbf{G}\in\mathbb{R}^{c\times q}$. Its $l$-rank approximation can be expressed as: $\mathbf{G}\approx\mathbf{C}\mathbf{Q}$, where $\mathbf{C}\in\mathbb{R}^{c\times l}$ and $\mathbf{Q}\in\mathbb{R}^{l\times q}$. When multiplying with a vector $\mathbf{b}\in\mathbb{R}^{q}$, we can first calculate the product $\mathbf{Q}\mathbf{b}$ and then multiply the resultant vector with $\mathbf{C}$. By such, the number of operations required can be reduced from O(cq){O}(cq) to O((c+q)l){O}((c+q)l). Since the function $\boldsymbol{\alpha}_{2}^{(t)}$ is asymptotically smooth, $\mathbf{G}$ can be well approximated by the $l$-rank approximation with $l\ll c$ and $l\ll q$, as long as $\mathbf{G}$ satisfies the admissibility condition [34, 35, 36]. Thus, the cost of the matrix-vector multiplication can be significantly reduced by the low-rank approximation. Given a configuration $\mathbf{X}$ of $N$ particles, we can employ the cluster-tree partitioning algorithm [39] to partition the associated matrix $\mathbf{M}^{(t,2)}$ into blocks such that each block satisfying the admissibility condition can be well approximated by a low-rank approximation. The resulting matrix, composed largely of low-rank blocks, is the so-called $\mathcal{H}$-matrix.
@@
In this work, we employ the adaptive cross approximation (ACA) [36] to find a low-rank approximation, thereby avoiding the computation of all entries of the original matrix. As a result, constructing $\mathcal{H}$-matrix requires ${O}(N\log N)$ operations, whereas the matrix-vector multiplication with an $\mathcal{H}$-matrix exhibits O(N){O}(N) complexity. More details about the cluster-tree block partitioning and ACA are provided as follows.